OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

rishavraj (rishavraj):

\[\frac{ y_2 - y_1 }{ x_2 - x_1 } = m\]

OpenStudy (campbell_st):

the slope intercept form of the line is y = mx + b you have been given the slope 3/5 so \[y = \frac{3}{5} x + b\] now substitute x = 2 and y = -4 to find b, and the equation of the line

OpenStudy (anonymous):

so 3/5(2)+(-4)?? @campbell_st

OpenStudy (campbell_st):

no \[-4 = \frac{3}{5}\times 2+ b\]

OpenStudy (anonymous):

so it'd be y+4=3/5(x-2)?? @campbell_st

OpenStudy (anonymous):

when you solve for b -4=3/5(2)+b -4=6/5+b b=-4-(6/5) =-26/5 so y=3/5x-26/5